If the selling price per unit is $250, total fixed expenses are $360,000, and the breakeven sales in dollars is $600,000, what is the variable expense per unit?

$100.00

$66.67

$150.00

$400.00
Answer #1

Breakeven sales=Fixed expense/Contribution margin ratio

Contribution margin ratio=(360,000/600,000)=0.6

Contribution margin ratio=Contribution margin/Sales

Contribution margin =(0.6*250)=$150

Contribution margin=Sales-Variable cost

Variable cost=(250-150)

=$100 per unit.